OpenStudy (anonymous):

5. which ordered pair is a solution of the equation y= -7x + 2? A. (1,2) B.(8,-54) C. (5,-35) D. (1,-7)

OpenStudy (anonymous):

@gorv

OpenStudy (gorv):

put each value

OpenStudy (gorv):

(1,2) y=-7*1+2=-5 y not equal to 2

OpenStudy (gorv):

so this is not our ans

OpenStudy (gorv):

(8,-54) y=-7*8+2=-56+2=-54

OpenStudy (gorv):

so here value in bracket and from equation are coming same ..so it is our ans

OpenStudy (anonymous):

so its B?

OpenStudy (anonymous):

got it thanks! can you help me with this one? -4=7/33x A. x= 132/7 B. x= -132/7 C. x=-7/132 D.x=-33/28

OpenStudy (gorv):

\[-4=\frac{ 7 }{ 33x }\]

OpenStudy (gorv):

multiply both side by x

OpenStudy (anonymous):

yes thats how its asked

OpenStudy (anonymous):

ok

OpenStudy (gorv):

\[-4*x=\frac{ 7*x }{ 338x }\]

OpenStudy (gorv):

now x on right side will get cancelled

OpenStudy (gorv):

sorry that is 33*x not 338x

OpenStudy (anonymous):

ok

OpenStudy (gorv):

-4x=7/33 x=7/(33*(-4))=-7/132 x=-7/132

OpenStudy (anonymous):

so its C, ok thanks! i have another one for you, -8=z/3 A. z=-8/3 B. z=-5 C. z=-24 D. z= 3/8

OpenStudy (gorv):

-8*3=3*z/3=z z=-24

OpenStudy (anonymous):

ok thanks! i still have some questions i need help on but im gonna take a lunch break.